At 2500km above the earth's surface the density of the atmosphere is about 1molecule/cm³

a) What is the mean free path predicted by pV = nRT.

b) What is its significance under these conditions? Assume a molecular diameter of 20 X 108cm


1
Expert's answer
2021-08-20T12:55:51-0400
"\lambda=\frac{1}{\sqrt{2}\pi d^2 n}\\
\lambda=\frac{1}{\sqrt{2}\pi (2\cdot10^{-10})^2(1\cdot10^{6})}=6\cdot10^{12}m"